OpenStudy (anonymous):

What are three raidcal expressions tgat simplify to 2xsqrt(3)?

OpenStudy (anonymous):

do you mean \[2 \times \sqrt{3}\] or \[2x \sqrt{3}\]

OpenStudy (anonymous):

I mean't 2x√3